OpenStudy (anonymous):

If 9^a *3^c = 177147, then 2a+c=? Thank you for your help! :)

OpenStudy (phi):

write 9 as \( 3^2\) so you have \[ 9^a \cdot 3^c = 177147 \\ \left(3^2\right)^a \cdot 3^c= 177147 \\ 3^{2a}\cdot 3^c= 177147 \\ 3^{2a+c}= 177147\] to find 2a+c, take the log of both sides. you will need a calculator.

OpenStudy (anonymous):

Wow, thank you very much! That's pretty complicated haha, but I got it thanks to you!